# Mixture-of-Experts with Gradient Conflict-Driven Subspace Topology Pruning for Emergent Modularity

### v1

@@ Abstract
Mixture-of-Experts (MoE) architectures achieve parameter efficiency through conditional computation, yet contemporary designs suffer from two fundamental limitations: structural parameter isolation that causes catastrophic forgetting, and instruction-overfitting that degrades performance in instruction-free scenarios. We propose CDSP-MoE (Conflict-Driven Subspace Pruning MoE), a framework that addresses these issues through a paradigm shift from isolated expert containers to dynamic expert instantiation within a shared physical subspace. Grounded in the Universal Weight Subspace Hypothesis, CDSP-MoE maintains a super-complete parameter backbone where logical experts are carved out via learnable topology masks. Unlike prior work that uses gradient conflict for token reassignment or optimization surgery, we leverage it as a structural supervisory signal: a Lagged Gradient Game penalizes interfering connections in the shared manifold, enabling the topology to spontaneously prune conflicting pathways and evolve interpretable modular structures. Experimental results demonstrate that CDSP-MoE achieves robust content-driven routing without human-defined task labels, maintaining semantic specialization even under strict blind inference protocols where explicit instructions are absent. Code is available at: https://github.com/konodiodaaaaa1/Conflict-Driven-Subspace-Pruning-Mixture-of-Experts
@@ 1 Introduction
The scaling of Large Language Models (LLMs) has been profoundly catalyzed by the Mixture-of-Experts (MoE) paradigm. By conditionally activating a subset of parameters, architectures like GShard [5] and Switch Transformer [2] have broken the dense scaling barrier. Recent advancements, such as DeepSeek-MoE [1], have further optimized this via fine-grained expert segmentation and shared-expert mechanisms, achieving remarkable efficiency.
@@
  • Memory Overwrite & Catastrophic Forgetting: The pursuit of load balancing often forces semantically unrelated tokens to update the same expert. Since experts are isolated containers without orthogonal subspaces, new gradient updates inadvertently overwrite previously learned knowledge, creating a “tug-of-war” dynamic that degrades performance in multi-task continuous learning.
@@
To address these limitations, we propose Mixture-of-Experts with Conflict-Driven Subspace Pruning (CDSP-MoE). Moving beyond the paradigm of isolated parameter containers, we ground our framework in the Universal Weight Subspace Hypothesis [3], which posits that diverse tasks can be solved by varying combinations of a shared, low-rank basis. In CDSP-MoE, logical experts are no longer disjoint tensors but dynamic instantiations carved out from a super-complete physical backbone via topology-aware masking.
The core of our methodology is the Lagged Gradient Game, a mechanism inspired by multi-task gradient surgery [8]. Instead of treating gradient conflict as optimization noise to be suppressed, we leverage it as a structural supervisory signal. By monitoring the cosine similarity of gradients between experts on the shared backbone, CDSP-MoE penalizes connections that cause interference. This forces the topology to spontaneously prune conflicting pathways, evolving a modular structure where experts specialize not by forced instruction IDs, but by minimizing physical interference.
@@
  • Robust Instruction-Free Performance: Extensive experiments show that CDSP-MoE achieves superior zero-shot routing accuracy and interpretability compared to DeepSeek-MoE and LoRA baselines under blind inference protocols.
@@ 2.1 Sparse Mixture-of-Experts
The concept of conditional computation was popularized by the Sparsely-Gated MoE [6], which introduced a noisy top-$k$ gating mechanism. Subsequent works like GShard [5] and Switch Transformer [2] focused on simplifying routing logic. More recently, DeepSeek-MoE [1] introduced fine-grained expert segmentation to enhance specialization.
@@
To address the load imbalance issue in top-$k$ routing, Expert Choice Routing [10] inverts the mechanism by allowing experts to select the top-$k$ tokens, thereby enforcing perfect capacity utilization. However, this relies on fixed capacity constraints which can drop critical information. CDSP-MoE differs fundamentally: we abandon forced load balancing entirely. Instead of optimizing for server capacity, we use conflict-driven pruning to optimize for semantic interference, allowing expert utilization to be governed solely by physical compatibility.
@@ 2.2 Subspace Learning and LoRA
Our work is grounded in the Universal Weight Subspace Hypothesis [3]. Low-Rank Adaptation (LoRA) [4] exploits this by training low-rank matrices.
@@
Our approach shares conceptual roots with AdaLoRA [9], which adaptively allocates parameter budgets by pruning singular values based on importance scores. While AdaLoRA optimizes the rank of static matrices for a specific dataset, CDSP-MoE operates on a shared physical manifold to perform instance-wise dynamic instantiation. This allows our model to reconfigure its effective topology on-the-fly for every token, rather than settling on a single static pruned structure.
@@ 2.3 Gradient Conflict in Multi-Task Learning
Gradient conflict arises when different tasks generate opposing parameter update directions, degrading multi-task optimization. Methods like PCGrad [8] address this through gradient surgery, projecting conflicting gradients to preserve task-specific directions during optimization.
Recent work has explored gradient conflict specifically within MoE architectures. STGC [7] identifies token-level gradient conflicts within each expert and introduces a conflict elimination loss to reassign conflicting tokens to alternative experts. While STGC improves routing decisions in vision-language models, it operates within the conventional MoE paradigm of largely isolated expert parameters: gradient conflict informs which tokens go to which expert, but the physical structure of experts remains architecturally fixed and independently parameterized.
@@ 3.3 Topology-Aware Dynamic Instantiation
To ensure stability while allowing structural plasticity, we adopt a Biased Weak-Connection initialization strategy:
$\mathbf{A}_{ij}\leftarrow\begin{cases}1.0&\text{if }i=j\quad\text{(Self-Preservation)}\\ 0.1+\epsilon&\text{if }i\neq j\quad\text{(Weak Cross-Link)}\end{cases}$ (3)
where $\epsilon\sim\mathcal{N}(0,0.01)$. This setup guarantees that experts possess functional autonomy at initialization while maintaining weak global connectivity for gradient exploration.
@@ 3.4 Forward Dynamics: Subspace Addressing and Computation
This mechanism effectively creates a "soft" relaxation of the hard selection process, enabling end-to-end learning of the topology.
@@ 3.6 Gradient Conflict Optimization
To avoid overhead, these gradients are detached and stored. In step $t+1$, we use the gradients $\mathbf{g}^{(t)}$ as "lagged signals" to optimize the topology.
@@
Structural Evolution and Gradient Flow.
The total objective function combines the task loss with the structural conflict penalty:
$\mathcal{L}_{total}=\mathcal{L}_{task}+\lambda_{conf}\sum_{i\neq j}\underbrace{\sigma(\mathbf{A}_{ij})\cdot\mathcal{C}_{ij}}_{\text{Conflict Loss}}+\lambda_{reg}\|\mathbf{A}\|_{1}$ (16)
Analysis of Gradient Flow: The optimization dynamics can be decomposed into two orthogonal pathways:
@@
    Subspace Learning ($\nabla_{\Theta}\mathcal{L}_{task}$): The task loss updates the values of the physical backbone $\Theta_{base}$, enabling feature extraction.
@@
    Topology Pruning ($\nabla_{\mathbf{A}}\mathcal{L}_{conf}$): The conflict loss generates gradients w.r.t the topology matrix $\mathbf{A}$. specifically:
    $\frac{\partial\mathcal{L}_{conf}}{\partial\mathbf{A}_{ij}}=\sigma^{\prime}(\mathbf{A}_{ij})\cdot\mathcal{C}_{ij}$ (17)
This derivation reveals the pruning mechanism: if expert $i$ has a strong connection to a subspace (high $\sigma(\mathbf{A}_{ij})$) that causes high conflict ($\mathcal{C}_{ij}$), the gradient $\frac{\partial\mathcal{L}_{conf}}{\partial\mathbf{A}_{ij}}$ becomes large and positive. Gradient descent then decreases $\mathbf{A}_{ij}$, effectively “cutting” the wire to that subspace. This forces the expert to migrate to a conflict-free region of the backbone.
@@ 3.7 Optimization Strategy
The higher rate for $\mathbf{A}$ allows the topology to adapt rapidly to the detected conflicts ("fast plasticity"), while the physical backbone consolidates knowledge slowly ("slow stability").
@@ 4 Experiments
Our experimental evaluation is designed not merely to chase marginal accuracy gains, but to verify the central hypothesis of this paper: that extitconflict-driven pruning can induce emergent modularity without human-defined semantic labels. We conduct controlled experiments in a heterogeneous multi-task environment to compare the structural evolution of CDSP-MoE against standard baselines.
@@ 4.1 Experimental Setup
We compare CDSP-MoE against a standard MoE baseline. To ensure a fair comparison focused on structural efficiency rather than capacity, we enforce a strict extitIso-Parameter Constraint:
@@
Models are trained using the AdamW optimizer. A key detail is the Two-Speed Learning Rate: the physical parameters are trained at $\eta=5e-3$, while the topology matrix $\mathbf{A}$ is evolved at $10\eta$ (5e25e-2). This differential rate is critical for allowing the structure to evolve faster than the accumulation of weight knowledge ("Evolutionary Tax"), ensuring that connections are pruned before they overfit to noise.
@@ 4.2 Evaluation Philosophy: From Identity to Content
Traditional MoE evaluations often prioritize peak accuracy on known tasks. However, this metric fails to capture whether the model has truly learned to route based on semantics or is simply memorizing Task ID mappings ("Identity Politics"). We propose a more rigorous evaluation protocol based on two core logics:
@@ 4.3.1 Experiment I: Emergent Modularity in CDSP-MoE
Figure 2 visualizes the sigmoid-activated topology matrix $\sigma(\mathbf{A})$ across training epochs.
@@
  • Pruning and Specialization (Epoch 9): Under the pressure of the conflict loss $\mathcal{L}_{conf}$ and L1 regularization, a clear "oligarchic" structure emerges. As shown in Figure 2(b), only a subset of experts (E5, E7, E2, E0) maintain high diagonal weights (>0.8>0.8), while others (E1, E3, E4, E6) are effectively marginalized. This confirms that CDSP-MoE spontaneously discovers the minimal effective parameter set.
(a) Epoch 0: Initialization
(b) Epoch 9: Oligarchic Convergence
Figure 2: Evolution of the topology matrix $\sigma(\mathbf{A})$ across training epochs.
@@
  • Stage 2: Semantic Convergence (Epoch 9). As the conflict game progresses, the model spontaneously realigns KMNIST with MNIST, routing both "symbolic" tasks to the E0/E5 cluster, while Fashion-MNIST migrates to a dedicated expert E7. This proves that the model successfully disentangles "symbolic" concepts from "object" concepts.
(a) Epoch 2: Complexity-Based Split
(b) Epoch 9: Semantic-Based Split
Figure 3: Comparison of expert utilization heatmaps between Epoch 2 and Epoch 9.
@@
Figure 4 presents the routing behavior when Task IDs are forcibly removed (Input None).
  • Semantic Robustness: The primary semantic boundary remains intact. The "Symbolic Cluster" (Real T0 and T1) is still dominantly routed to E0 and E5, while the "Object Task" (Real T2) maintains its distinct preference for E7.
@@
Figure 4: Routing distribution of CDSP-MoE under blind inference (Task ID = None).
@@ 4.3.2 Experiment II: Instruction Overfitting in Standard Baselines
Figure 5 illustrates the routing distribution of the Baseline at Epoch 9 with explicit Task IDs.
@@
  • The Lookup Table Mechanism: While this appears effective, it reveals a reliance on "shortcut learning." The router learns a simple linear mapping from $ID\to Expert$, bypassing the need to analyze the visual features of the input.
Figure 5: Routing distribution of the Standard MoE Baseline in the training state.
@@
The fragility of the baseline is exposed in the blind inference test (Figure 6), where Task IDs are removed.
@@
  • Loss of Specialized Knowledge: Expert 2, the "Clothing Expert" during training, is abandoned. The router defaults to experts with the highest global frequency bias (E4 and E6), demonstrating a complete lack of content-driven decision making.
Figure 6: Routing distribution of the Standard MoE Baseline under blind inference.

### v3

@@ Abstract
Mixture-of-Experts (MoE) architectures achieve parameter efficiency through conditional computation, yet contemporary designs suffer from structural parameter isolation. We propose CDSP-MoE (Conflict-Driven Subspace Pruning MoE), a framework that aims to alleviate these structural bottlenecks through a paradigm shift from isolated expert containers to dynamic expert instantiation within a shared physical subspace. Grounded in the Universal Weight Subspace Hypothesis, CDSP-MoE maintains a super-complete parameter backbone where logical experts are carved out via learnable topology masks. Unlike prior work that uses gradient conflict for token reassignment or optimization surgery, we leverage it as a structural supervisory signal: a Lagged Gradient Game penalizes interfering connections in the shared manifold, enabling the topology to spontaneously prune conflicting pathways and evolve interpretable modular structures. Experimental results demonstrate that CDSP-MoE achieves robust content-driven routing without human-defined task labels, maintaining semantic specialization even under strict blind inference protocols where explicit instructions are absent.
MoE, Gradient Conflict, Subspace Pruning, ICML
@@ 1 Introduction
The scaling of Large Language Models (LLMs) has been profoundly catalyzed by the Mixture-of-Experts (MoE) paradigm. By conditionally activating a subset of parameters, architectures like GShard (Lepikhin et al., 2020) and Switch Transformer (Fedus et al., 2021) have broken the dense scaling barrier. Recent advancements, such as DeepSeek-MoE (Dai et al., 2024), have further optimized this via fine-grained expert segmentation and shared-expert mechanisms, achieving remarkable efficiency.
@@
  • Memory Overwrite & Catastrophic Forgetting: The pursuit of load balancing often forces semantically unrelated tokens to update the same expert. Since experts are isolated containers without orthogonal subspaces, new gradient updates inadvertently overwrite previously learned knowledge, creating a “tug-of-war” dynamic that is associated with performance in multi-task continuous learning.
@@
To explore a structural route toward mitigating for these limitations, we propose Mixture-of-Experts with Conflict-Driven Subspace Pruning (CDSP-MoE). Moving beyond the paradigm of isolated parameter containers, we ground our framework in the Universal Weight Subspace Hypothesis (Kaushik et al., 2025), which posits that diverse tasks can be solved by varying combinations of a shared, low-rank basis. In CDSP-MoE, logical experts are no longer disjoint tensors but dynamic instantiations carved out from a super-complete physical backbone via topology-aware masking.
The core of our methodology is the Lagged Gradient Game, a mechanism inspired by multi-task gradient surgery (Yu et al., 2020). Instead of treating gradient conflict as optimization noise to be suppressed, we leverage it as a structural supervisory signal. By monitoring the cosine similarity of gradients between experts on the shared backbone, CDSP-MoE penalizes connections that cause interference. This forces the topology to spontaneously prune conflicting pathways, evolving a modular structure where experts specialize not by forced instruction IDs, but by minimizing physical interference.
While the motivation stems from the challenges in scaling LLMs, CDSP-MoE proposes a fundamental shift in neural topology evolution. To rigorously isolate and visualize the emergent modularity without the interference of complex language semantics, we evaluate our framework on a heterogeneous multi-task benchmark (MNIST/Fashion-MNIST), serving as a theoretical proof-of-concept for the underlying mechanism.
@@
  • Robust Instruction-Free Performance: Extensive experiments show that CDSP-MoE achieves superior zero-shot routing accuracy and interpretability compared to Standard MoE baselines under blind inference protocols.
@@ 2.1 Sparse Mixture-of-Experts
The concept of conditional computation was popularized by the Sparsely-Gated MoE (Shazeer et al., 2017), which introduced a noisy top-$k$ gating mechanism. Subsequent works like GShard (Lepikhin et al., 2020) and Switch Transformer (Fedus et al., 2021) focused on simplifying routing logic. More recently, DeepSeek-MoE (Dai et al., 2024) introduced fine-grained expert segmentation to enhance specialization.
@@
To address the load imbalance issue in top-$k$ routing, Expert Choice Routing (Zhou et al., 2022) inverts the mechanism by allowing experts to select the top-$k$ tokens, thereby enforcing perfect capacity utilization. However, this relies on fixed capacity constraints which can drop critical information. CDSP-MoE differs fundamentally: we abandon forced load balancing entirely. Instead of optimizing for server capacity, we use conflict-driven pruning to optimize for semantic interference, allowing expert utilization to be governed solely by physical compatibility.
@@ 2.2 Subspace Learning and LoRA
Our work is grounded in the Universal Weight Subspace Hypothesis (Kaushik et al., 2025). Low-Rank Adaptation (LoRA) (Hu et al., 2022) exploits this by training low-rank matrices.
@@
Our approach shares conceptual roots with AdaLoRA (Zhang et al., 2023), which adaptively allocates parameter budgets by pruning singular values based on importance scores. While AdaLoRA optimizes the rank of static matrices for a specific dataset, CDSP-MoE operates on a shared physical manifold to perform instance-wise dynamic instantiation. This allows our model to reconfigure its effective topology on-the-fly for every token, rather than settling on a single static pruned structure.
@@ 2.3 Gradient Conflict in Multi-Task Learning
Gradient conflict arises when different tasks generate opposing parameter update directions, degrading multi-task optimization. Methods like PCGrad (Yu et al., 2020) address this through gradient surgery, projecting conflicting gradients to preserve task-specific directions during optimization. While these methods effectively stabilize optimization within a fixed architecture, CDSP-MoE extends this perspective to the structural level, utilizing conflict signals to physically decouple pathways and induce emergent modularity.
Recent work has explored gradient conflict specifically within MoE architectures. STGC (Yang et al., 2025) identifies token-level gradient conflicts within each expert and introduces a conflict elimination loss to reassign conflicting tokens to alternative experts. While STGC improves routing decisions in vision-language models, it operates within the conventional MoE paradigm of largely isolated expert parameters: gradient conflict informs which tokens go to which expert, but the physical structure of experts remains architecturally fixed and independently parameterized.
@@ 3.3 Topology-Aware Dynamic Instantiation
Instead of a sparse initialization which may hinder early exploration, we adopt a Maximum Entropy initialization strategy. We initialize the topology logits $\mathbf{A}$ near zero, placing the system in a highly plastic ”semi-connected” state ($P\approx 0.5$):
$\mathbf{A}_{ij}\leftarrow\begin{cases}4.0&\text{if }i=j\quad\text{(Strong Self-Preservation)}\\ 0+\epsilon&\text{if }i\neq j\quad\text{(Maximum Plasticity)}\end{cases}$ (3)
where $\epsilon\sim\mathcal{N}(0,0.02)$. This unbiased starting point ensures that gradient flow is not suppressed by saturation regions of activation functions, allowing the Conflict and Synergy signals to sculpt the topology effectively from the very first step.
@@ 3.4 Forward Dynamics: Subspace Addressing and Computation
This mechanism creates a ”soft” relaxation of the hard selection process. While $m_{i}$ is a scalar, it serves as a sufficient coarse-grained gatekeeper: if the subspace as a whole is effective, the scalar feedback reinforces the corresponding topology weights, enabling end-to-end learning without the instability of complex vector-wise estimators.
@@ 3.6 Gradient Conflict Optimization
To avoid overhead, these gradients are detached and stored. In step $t+1$, we use the gradients $\mathbf{g}^{(t)}$ as ”lagged signals” to optimize the topology.
@@
Structural Evolution and Objective Function.
The total objective function operates directly on the topology logits to avoid gradient vanishing issues:
$\mathcal{L}_{total}=\mathcal{L}_{task}+\lambda_{conf}\sum_{i\neq j}\underbrace{\mathbf{A}_{ij}\cdot\mathcal{C}_{ij}}_{\text{Direct Conflict Penalty}}+\lambda_{reg}\|\mathbf{A}\|_{1}$ (16)
By removing the sigmoid scaling factor from the penalty term, we ensure that the pruning force remains proportional to the conflict magnitude $\mathcal{C}_{ij}$, regardless of the current connection strength.
Analysis of Gradient Flow: The optimization dynamics benefits from the linearity of the logit-space penalty:
@@
    Subspace Learning ($\nabla_{\Theta}\mathcal{L}_{task}$): The task loss updates the values of the physical backbone $\Theta_{base}$.
@@
    Topology Pruning ($\nabla_{\mathbf{A}}\mathcal{L}_{conf}$): The conflict loss generates direct gradients w.r.t the topology logits $\mathbf{A}$:
    $\frac{\partial\mathcal{L}_{conf}}{\partial\mathbf{A}_{ij}}=\mathcal{C}_{ij}$ (17)
Mechanism Clarification: Unlike sigmoid-gated penalties where gradients vanish when connections are strong ($\sigma\approx 1$) or weak ($\sigma\approx 0$), our formulation applies a constant pruning pressure. If conflict exists ($\mathcal{C}_{ij}>0$), gradient descent linearly decreases $\mathbf{A}_{ij}$, rapidly pushing the connection towards negative values (disconnection) without saturation delays.
The Role of Regularization.
The $L_{1}$ regularization on logits ($\|\mathbf{A}\|_{1}$) plays a distinct role from standard sparsity induction. Since $\mathbf{A}_{ij}=0$ corresponds to a probability of $0.5$, minimizing $\|\mathbf{A}\|_{1}$ acts as a centering force, pulling parameters towards the unbiased initialization state. This prevents inactive connections from drifting into deep negative saturation (dead zones) and maintains their plasticity, allowing them to be ”resurrected” if task demands change.
@@ 3.7 Optimization Strategy
The higher rate for $\mathbf{A}$ allows the topology to adapt rapidly to the detected conflicts (”fast plasticity”), while the physical backbone consolidates knowledge slowly (”slow stability”). Crucially, this timescale separation ensures that the physical landscape remains relatively stable between steps, validating the use of lagged gradients $\mathbf{g}^{(t)}$ as a reliable approximation for structural interference at step $t+1$.
@@ 4 Experiments
Our experimental evaluation is designed not merely to chase marginal accuracy gains, but to verify the central hypothesis of this paper: that conflict-driven pruning can induce emergent modularity without human-defined semantic labels. We conduct controlled experiments in a heterogeneous multi-task environment to compare the structural evolution of CDSP-MoE against standard baselines.
@@ 4.1 Experimental Setup
We compare CDSP-MoE against a standard MoE baseline. To ensure a fair comparison focused on structural efficiency rather than capacity, we enforce a strict Iso-Parameter Constraint:
@@
Models are trained using the AdamW optimizer. A key detail is the Two-Speed Learning Rate: the physical parameters are trained at $\eta=5e-3$, while the topology matrix $\mathbf{A}$ is evolved at $10\eta$ (5e25e-2). This differential rate is critical for allowing the structure to evolve faster than the accumulation of weight knowledge (”Evolutionary Tax”), ensuring that connections are pruned before they overfit to noise.
@@ 4.2 Evaluation Philosophy: From Identity to Content
Traditional MoE evaluations often prioritize peak accuracy on known tasks. However, this metric fails to capture whether the model has truly learned to route based on semantics or is simply memorizing Task ID mappings (”Identity Politics”). We propose a more rigorous evaluation protocol based on two core logics:
@@ 4.3 Results and Analysis
(a) Topology (Epoch 0)
(b) Topology (Epoch 9)
(c) Routing (Epoch 2)
(d) Routing (Epoch 9)
Figure 2: Structural Evolution of CDSP-MoE. (a-b) The topology matrix $\sigma(\mathbf{A})$ transitions from uniform initialization to sparse oligarchy. (c-d) Routing heatmaps show the shift from complexity-based grouping to semantic specialization.
@@ 4.3.1 Experiment I: Emergent Modularity in CDSP-MoE
Figure A4 visualizes the sigmoid-activated topology matrix $\sigma(\mathbf{A})$ across training epochs.
@@
  • Pruning and Specialization (Epoch 9): Under the pressure of the conflict loss $\mathcal{L}_{conf}$ and L1 regularization, a clear ”oligarchic” structure emerges. As shown in Figure A4, only a subset of experts (E5, E7, E2, E0) maintain high diagonal weights (>0.8>0.8), while others (E1, E3, E4, E6) are effectively marginalized. This confirms that CDSP-MoE spontaneously discovers the minimal effective parameter set.
@@
  • Stage 2: Semantic Convergence (Epoch 9). As the conflict game progresses, the model spontaneously realigns KMNIST with MNIST, routing both ”symbolic” tasks to the E0/E5 cluster, while Fashion-MNIST migrates to a dedicated expert E7. This suggests that the model successfully disentangles ”symbolic” concepts from ”object” concepts.
(a) Baseline (Train)
(b) Baseline (Blind)
(c) Pure Blind (Train)
(d) CDSP (Blind)
Figure 3: Comparative Routing Analysis. (a) Standard Baseline overfits to Task IDs. (b) Baseline collapses under blind inference. (c) Pure Blind Baseline fails to isolate semantics (entanglement). (d) CDSP maintains robust semantic modularity even without Task IDs.
@@
Figure 3(d) presents the routing behavior when Task IDs are forcibly removed (Input None).
  • Semantic Robustness: The primary semantic boundary remains intact. The ”Symbolic Cluster” (Real T0 and T1) is still dominantly routed to E0 and E5, while the ”Object Task” (Real T2) maintains its distinct preference for E7.
@@ 4.3.2 Experiment II: Instruction Overfitting in Standard Baselines
Figure 3(a) illustrates the routing distribution of the Baseline at Epoch 9 with explicit Task IDs.
@@
  • The Lookup Table Mechanism: While this appears effective, it reveals a reliance on ”shortcut learning.” The router learns a simple linear mapping from $ID\to Expert$, bypassing the need to analyze the visual features of the input.
@@
The fragility of the baseline is exposed in the blind inference test (Figure 3(b)), where Task IDs are removed.
@@
  • Loss of Specialized Knowledge: Expert 2, the ”Clothing Expert” during training, is abandoned. The router defaults to experts with the highest global frequency bias (E4 and E6), demonstrating a complete lack of content-driven decision making.